The concentrations of reactants and products for a
Answer #1

B)

CO + NH3 = HCONH2

K = 0.77

CO = 1

NH3 = 2

HCONH2 = 0

in equilibrium

CO = 1 - x

NH3 = 2 - x

HCONH2 = 0 + x

then

K = HCONH2 /CO*NH3

0.77 = (x)/(2-x)(1-x)

1.29870129x = 2-3x+x^2

x^2 - (1.29870129 +3)x +2 = 0

x = 0.53

since

HCONH2 = 0 + x = 0.53M